, The solutions of the inequality IIxI-1I <=1/2 are ?



1
Expert's answer
2021-09-02T12:30:49-0400

"\lvert \lvert x \rvert - 1\rvert \leq \cfrac {1}{2}"


There are 3 cases:


1) "\lvert x \rvert - 1 = 0"

2) "\lvert x \rvert - 1 > 0"

3) "\lvert x \rvert - 1 < 0"


1) "\lvert x \rvert - 1 = 0"

"\lvert x \rvert = 1 \\
x \in \lbrace -1; 1\rbrace" - satisfies the inequality.


2) "\lvert x \rvert - 1 > 0 :"


"0 < \lvert x \rvert - 1 \leq \cfrac{1}{2} \\
1 < \lvert x \rvert \leq \cfrac{3}{2}"


Answer: "x \in \Big [ - \cfrac {3}{2} ; -1 \Big \} \bigcup \Big \{ 1; \cfrac{3}{2} \Big ]"


3) "\lvert x \rvert - 1 < 0 :"


"-\cfrac{1}{2} \leq \lvert x \rvert - 1 < 0 \\
\cfrac{1}{2} \leq \lvert x \rvert < 1"


Answer: "x \in \Big \{ - 1 ; - \cfrac{1}{2} \Big ] \bigcup \Big [ \cfrac{1}{2} ; 1 \Big \}"



Final answer: "x \in \Big [ - \cfrac{3}{2}; - \cfrac{1}{2} \Big ] \bigcup \Big [ \cfrac{1}{2} ; \cfrac {3}{2} \Big ]"